HIGH RESOLUTION LASER SPECTROSCOPY OF THE [20.7]^{3}Φ_{4} - X^{3}Φ_{4} AND [21.3]^{3}Φ_{4} - X^{3}Φ_{4} SYSTEMS OF COBALT MONOCHLORIDE.

YING TENG and \underline{A. G. ADAM}, Department of Chemistry, University of New Brunswick, Fredericton, NB, Canada E3B 6E2; J. R. D. PEERS, Department of Chemistry, University of British Columbia, Vancouver, BC, Canada V6T 1Z1.

Low and high resolution spectra of cobalt monochloride have been observed in the 455 - 485 nm region using a laser ablation/molecular beam spectrometer. The CoCl molecules were produced by laser vaporization of a Co rod followed by reaction with CCl_{4} inside a pulsed supersonic jet source. A number of bands were seen allowing vibrational assignments to be made to transitions between the ground state and two excited electronic states with 0-0 origins near 20680.5 and 21257.3 cm^{−1}. The spectra of three vibrational bands, the 0-0 and 1-0 bands of the [20.7]^{3}Φ_{4} - X^{3}Φ_{4} transition and the 0-0 band of the [21.3]^{3}Φ_{4} - ^{3}Φ_{4} transition, were recorded at high resolution with a Coherent 699-29 ring dye laser. A rotational and hyperfine analysis has been performed on these bands giving an accurate set of molecular constants for the ground and excited states. The assignment of a ^{3}Φ state as the ground state is consistent with recent experimental and theoretical predictions for CoH and CoF.